I am given the concentration of a substance in a solution (2.7 mg/100ml) and fhe molecular weight of the substance (180.158g/mol) how do i find the quantity in g of substance?
Answer #1

From a given value: concentration (C) = 2.7 mg/100mL = (2.7*10^-3 g) / 100 mL = 2.7*10^-5 g/mL

Thus in one mL of solution we have 2.7*10^-5 g of solute.

We do not need molecular weight to calculate the amount of substance in 1 mL.

If you have given how much mL of the solution then we can determine the total amount of solute in g. Otherwise, we can determine only the number of g of solute in one mL of solution.

Please let me know in the comment section if you have been given with total volume of the solution then it is possible to calculate total g of solute in that solution. Accordingly, I will update my answer.